find the exact value of tan315 using a sum or difference identity.
Answer by lwsshak3(6505)

find the exact value of tan315 using a sum or difference identity.
**
tan(315)=tan(360-45)=(tan 360-tan 45)/(1+tan 360*tan 45)
tan 360º=0
tan 45º=1
(tan 360-tan 45)/(1+tan 360*tan 45)
(0-1)/(1+0*1)=-1/1=-1
ans:
tan 315º=-1